# Approximating the coefficients of the Bessel functions

### v4

@@ 1.8. Equivalence conditions for non-exponential holomorphic functions
  • In Theorems˜8.21 and 8.26, we set $F_{N}(x_{1},\ldots,x_{N})=\sum_{\lambda\in\Gamma_{\text{even};\,N}}(c_{\lambda}(N)+ed_{\lambda}(N))p_{\lambda}$, replace $\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}\right)$ with $\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}$, and replace $\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}d_{\gamma}p_{\gamma}\right)$ with $\sum_{\gamma\in\Gamma_{\text{even}}}d_{\gamma}p_{\gamma}$.
@@ 1.9. Paper organization
In Section˜2, we define the Dunkl operators and notation regarding partitions and noncrossing partitions. In Section˜3, we introduce the setting of applying a graded ring of operators to a graded vector field and in Section˜4, we define the notion of an invertible graded ring of operators. Afterwards, in Section˜5, we discuss a representation of invertible graded rings of operators as sequences of invertible matrices and connect the framework introduced in Section˜3 to the Dunkl operators. In Sections 6, 7, and 8, we discuss the leading order terms of the Dunkl bilinear form and prove Theorem˜1.1 for the $A^{N-1}$, $BC^{N}$, and $D^{N}$ root systems, respectively. Following this, in Section˜9, we determine the asymptotics of the coefficients of the terms of the Bessel functions that are homogeneous with a fixed degree. In Section˜10, we discuss applications of the results of this paper and in Section˜11, we present combinatorial expressions for the Dunkl bilinear form. In Section˜12, we analyze a special case of the leading order terms of the Dunkl bilinear form for the $A^{N-1}$ and $BC^{N}$ root systems.
@@ 2.1. Dunkl operators
The following result contains some properties about the eigenfunction $E_{a}^{\mathcal{R}(\theta)}(x)$.
@@ 2.2. Partitions
Additionally, define the unique coefficients $c_{\epsilon\lambda}$ for $\epsilon,\lambda\in\Gamma$ such that
@@
When we write $M_{\epsilon}=\sum_{\lambda\in\Gamma}c_{\epsilon\lambda}p_{\lambda}$, we mean that $M_{\epsilon}(x)=\sum_{\lambda\in\Gamma}c_{\epsilon\lambda}p_{\lambda}(x)$, where $x=(x_{i})_{i\geq 1}$ consists of an infinite number of variables. We could equivalently state that $M_{\epsilon}(x_{1},\ldots,x_{N})=\sum_{\lambda\in\Gamma}c_{\epsilon\lambda}p_{\lambda}(x_{1},\ldots,x_{N})$ for all $N\geq 1$.
@@ 2.3. Non-crossing partitions
Suppose $k\geq 1$. Let $NC(k)$ denote the set of noncrossing partitions of $[k]$. Recall that a partition $B_{1}\sqcup\cdots\sqcup B_{m}$ of $[k]$ is noncrossing if there does not exist distinct $i,j\in[m]$, $a,c\in B_{i}$, and $b,d\in B_{j}$ such that $a<b<c<d$. Furthermore, for $k\geq 1$, define $NC^{\text{even}}(k)$ to be the set of $\pi\in NC(k)$ such that each block of $\pi$ has even size.
@@
[…]
Furthermore, define $NC^{\prime}(k-1)$ to be the set of $\pi$ such that $m\geq 2$ and the minimal element of $B_{2}$ is $2$.
@@ 3. An introduction to applying a graded ring of operators to a graded vector field
The paper [dunkl_singular_poly] considers the applications of operators to a graded vector field $V\triangleq\oplus_{i\geq 0}V_{i}$, where the application of an operator to an element of $V_{i}$ outputs an element of $V_{i-1}$ for $i\geq 1$. When $V=\mathbb{C}[x_{1},\ldots,x_{N}]$ and $V_{i}$ is the set of elements of $V$ that are homogeneous of degree $i$ for all $i\geq 0$, we have that $\partial_{i}$, $\mathcal{D}(\mathcal{R}(\theta))_{i}$, and $d_{i}$ for $i\in[N]$ are examples of such an operator. In this section, we introduce a similar setting that is motivated by the Dunkl bilinear form $\cdot,\cdot]_{\mathcal{R}(\theta)}$ and we continue to discuss this setting in Sections˜4 and 5. See Example˜5.19 for the application of the framework that we introduce to the context of Dunkl operators.
@@ 4. Invertible graded rings of operators
The key contribution that we discuss in this section while proving Theorem˜4.5 is the analysis of the matrices $\{M^{i;L}\}_{i\geq 0}$. To the best of our knowledge, the specific method we use has not appeared previously. It allows for straightforward proofs that are applicable in a general setting.
@@
    For all $i\geq 0$, there does not exist nonzero $f\in R_{i}$ such that $L(g)f=0$ for all $g\in R_{i}$.
@@
    For all $i\geq 0$, there does not exist nonzero $g\in R_{i}$ such that $L(g)f=0$ for all $f\in R_{i}$.
@@
    The equation $\sum_{j=1}^{|A_{i}|}(L(b_{ij})f)a_{ij}=0$ has no nonzero solutions $f\in R_{i}$ for all $i\geq 0$.
@@ 5.2. Equivariance with respect to a group action and examples
It does not seem to be generally true that the invertibility of $L_{(R,V)}$ implies the invertibility of $L$. However, in the paper [dunkl_singular_poly], this is shown to be true for Dunkl operators. We repeat the argument from Remark 4.2 of that paper in a more general setting to prove the following result.
@@
For all $h\in H$, $hg^{\prime}=\prod_{h^{\prime}\in H}hh^{\prime}g=g^{\prime}$ by the commutativity of $S$, so $g^{\prime}\in R_{M}$, where $M$ is the positive integer such that $g^{\prime}\in S_{M}$.
@@
Suppose $f\in V_{M}$. Then, it is evident that $L(g^{\prime})f=0$ after applying $L(h^{\prime}g)$ for some sequence of |H|1|H|-1 distinct $h^{\prime}\in H\backslash\{1\}$. Thus, $L_{(R,V)}$ is not invertible, which is a contradiction. ∎
Remark 5.12.
In the previous result, we do not assume that $hS_{i}\subset S_{i}$ for all $h\in H$ and $i\geq 1$. In particular, we do not necessarily require that $M=|H|i$, although $M\geq 1$ because $g\in S_{i}$ and $i\geq 1$.
@@
Suppose $N\geq 1$, $S=W=K[x_{1},\ldots,x_{N}]$, and that $S_{i}=W_{i}$ is the group of polynomials over $K$ that are homogeneous of degree $i$ for $i\geq 0$. From Lemma˜5.15, any $L\in\mathcal{T}$ is uniquely determined by $L(x_{i})$ for $1\leq i\leq N$. Let $H$ be a finite group which acts on $K^{N}$ such that for all $h\in H$ and $f\in S_{1}$, $\{x\mapsto f(hx)\}\in S_{1}$.
@@
Assume that $L\in\mathcal{H}$ is invertible and that $\text{char}(K)$ does not divide $|H|$. Then, all conclusions of Corollary˜5.17 are true (this includes the conclusions over both $(R,V)$ and $(S,W)$).
@@
We describe the framework that we have developed in the context of Dunkl operators and prove a few well-known results. Consider the setting of Example˜5.14, but set $K=\mathbb{C}$ and $H$ to be $H(\mathcal{R})$, where $\mathcal{R}\subset\mathbb{R}^{N}$ is a finite root system.
Suppose $u_{i}\in\mathbb{R}^{N}$ for $1\leq i\leq N$. Then, we can define $L\in\mathcal{T}$ by $L(x_{i})\triangleq\mathcal{D}_{u_{i}}(\mathcal{R}(\theta))$ after using Lemma˜5.15 and the commutativity of the Dunkl operators. As a special case, we have that $\mathcal{D}(\mathcal{R}(\theta))\in\mathcal{T}$, where we recall that $\mathcal{D}(\mathcal{R}(\theta))(x_{i})=\mathcal{D}_{i}(\mathcal{R}(\theta))$ for $i\in[N]$. For $i\geq 0$, the matrix $M^{i;\,\mathcal{D}(\mathcal{R}(\theta))}$ has entry $(r,s)$ equal to $r,s]_{\mathcal{R}(\theta)}$ for $r\in D_{i}$ and $s\in C_{i}$.
@@
Observe that $E_{a}^{\mathcal{R}(\theta)}(x)$ is the $\Psi$-eigenvector of $\mathcal{D}(\mathcal{R}(\theta))$ and $J_{a}^{\mathcal{R}(\theta)}(x)$ is the $\Psi$-eigenvector of $\mathcal{D}_{H}(\mathcal{R}(\theta))$ for $\Psi:r\mapsto r(a)$, where $a\in\mathbb{C}^{N}$ is fixed. Then, part (B) of Corollary˜5.17 implies that these functions are unique eigenvectors, which is included in the statements of Theorems˜2.5 and 2.8.
We do not consider intertwining operators in the Dunkl setting in this paper. However, we mention that the intertwining operator for $(\mathcal{D}(\mathcal{R}(\theta)),\partial)$ is well-studied, where $\partial\in\mathcal{T}$ satisfies $\partial(x_{i})=\partial_{i}$ for $i\in[N]$; equivalently, $\partial\triangleq\mathcal{D}(\mathcal{R}(0))$. As an example of an application of Theorem˜4.5, the theorem implies the well-known result that $(\mathcal{D}(\mathcal{R}(\theta)),\partial)$ intertwines with a unique operator if and only if $\mathcal{D}(\mathcal{R}(\theta))$ is invertible. As explained in [rosler_positivity], in terms of Theorem˜1.4, the intertwiner is given by
$Vf(a)\triangleq\int_{\mathbb{R}^{N}}f(\epsilon)d\mu_{a}(\epsilon)$
for $f\in\mathbb{C}[x_{1},\ldots,x_{N}]$ and $a\in\mathbb{R}^{N}$.
@@
For $\theta\geq 0$ and even $i\geq 0$, it is straightforward to compute the decomposition of $M^{i;\,\mathcal{D}_{H}(D^{N}(\theta))}$ when we restrict the rows and columns to $p_{\lambda}$ for $\lambda\in\Gamma_{\text{even};\,N}[i]$ by using the decomposition of $M^{i;\,\mathcal{D}_{H}(BC^{N}(\theta,0))}$. An interesting question is, what is the decomposition when we restrict the rows and columns to $ep_{\lambda}$ for $\lambda\in\Gamma_{\text{even};\,N}[i-N]$, where $i\geq N$ has the same parity as $N$? Furthermore, another interesting question which generalizes the previous question is, what are the decompositions of $M^{i;\,\mathcal{D}(BC^{N}(\theta_{0},\theta_{1}))}$ for $i\geq 0$? Recall that $M^{i;\,\mathcal{D}(D^{N}(\theta))}=M^{i;\,\mathcal{D}(BC^{N}(\theta,0))}$.
@@ 6. Leading order terms for the type A Dunkl bilinear form
6.2. Proof of part (A) of Theorem˜1.1
We are now prepared to prove part (A) of Theorem˜1.1.
@@
The next theorem generalizes part (A) of Theorem˜1.1. As we mentioned previously, we focus on the less general version stated in Corollary˜6.11 because it is more relevant for analyzing the convergence of Bessel functions.
@@
6.3. Leading order terms of the type A Dunkl bilinear form in the $\theta N\rightarrow c\in\mathbb{C}$ regime
Next, we compute the leading order terms of the matrix $([p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)})_{\lambda,\nu\in\Gamma}$ in the regime $\theta N\rightarrow c\in\mathbb{C}$.

### v5

@@ 1.10. Equivalence conditions for non-exponential holomorphic functions
  • In Theorems˜8.9 and 8.11, we set $F_{N}(x_{1},\ldots,x_{N})=\sum_{\lambda\in\Gamma_{\text{even};\,N}}(c_{\lambda}(N)+ed_{\lambda}(N))p_{\lambda}$, replace $\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}\right)$ with $\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}$, and replace $\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}d_{\gamma}p_{\gamma}\right)$ with $\sum_{\gamma\in\Gamma_{\text{even}}}d_{\gamma}p_{\gamma}$.
@@ 1.11. Paper organization
In Section˜2, we define the Dunkl operators and notation regarding partitions and noncrossing partitions. In Section˜3, we introduce the setting of applying a graded ring of operators to a graded vector field and in Section˜4, we define the notion of an invertible graded ring of operators. Afterwards, in Section˜5, we discuss a representation of invertible graded rings of operators as sequences of invertible matrices and connect the framework introduced in Section˜3 to the Dunkl operators. In Sections 6, 7, and 8, we discuss the leading order terms of the Dunkl bilinear form and prove Theorem˜1.2 for the $A^{N-1}$, $BC^{N}$, and $D^{N}$ root systems, respectively. Following this, in Section˜9, we determine the asymptotics of the coefficients of the terms of the Bessel functions that are homogeneous with a fixed degree. In Section˜10, we discuss applications of the results of this paper. Following this, in Section˜11, we present combinatorial expressions for the Dunkl bilinear form.
@@ 2.1. Dunkl operators
The following lemma contains some properties about the eigenfunction $E_{a}^{\mathcal{R}(\theta)}(x)$. See Lemma˜10.4 for an upper bound on the symmetric eigenfunction that is similar to part (e) of the lemma.
@@ 2.2. Partitions
Additionally, define the unique coefficients $c_{\epsilon\lambda}\in\mathbb{Q}$ for $\epsilon,\lambda\in\Gamma$ such that
@@
We define $c_{\lambda\epsilon}^{\text{inv}}$ for $\epsilon,\lambda\in\Gamma$ such that
$p_{\lambda}=\sum_{\epsilon\in\Gamma}c_{\lambda\epsilon}^{\text{inv}}M_{\epsilon}.$
@@
When we write $M_{\epsilon}=\sum_{\lambda\in\Gamma}c_{\epsilon\lambda}p_{\lambda}$, we mean that $M_{\epsilon}(x)=\sum_{\lambda\in\Gamma}c_{\epsilon\lambda}p_{\lambda}(x)$, where $x=(x_{i})_{i\geq 1}$ consists of an infinite number of variables. We could equivalently state that $M_{\epsilon}(x_{1},\ldots,x_{N})=\sum_{\lambda\in\Gamma}c_{\epsilon\lambda}p_{\lambda}(x_{1},\ldots,x_{N})$ for all $N\geq 1$. This is similarly the case when we write $p_{\lambda}=\sum_{\epsilon\in\Gamma}c_{\lambda\epsilon}^{\text{inv}}M_{\epsilon}$. It is clear that $c_{\lambda\lambda}=\frac{1}{\pi(\lambda)}$ and $c_{\lambda\lambda}^{\text{inv}}=\pi(\lambda)$.
@@ 2.3. Non-crossing partitions
Suppose $k\geq 1$. Let $NC(k)$ denote the set of noncrossing partitions of $[k]$. Recall that a partition $B_{1}\sqcup\cdots\sqcup B_{m}$ of $[k]$ is noncrossing if there does not exist distinct $i,j\in[m]$, $a,c\in B_{i}$, and $b,d\in B_{j}$ such that $a<b<c<d$. Furthermore, for $k\geq 1$, define $NC^{\text{even}}(2k)$ to be the set of $\pi\in NC(2k)$ such that each block of $\pi$ has even size.
@@
2.4. Additional notation
For a measure $\mu$ over $\mathbb{R}^{N}$ and measurable $f:\mathbb{R}^{N}\rightarrow\mathbb{R}$, we let $\mu(f)$ denote the pushforward of $\mu$ with respect to the map $f$. For example, for a point mass $\delta_{a}$ located at $a\in\mathbb{R}^{N}$, $\delta_{a}(f)$ denotes the point mass $\delta_{f(a)}$.
@@ 3. An introduction to applying a graded ring of operators to a graded vector field
The paper [dunkl_singular_poly] considers the applications of operators to a graded vector field $V\triangleq\oplus_{i\geq 0}V_{i}$, where the application of an operator to an element of $V_{i}$ outputs an element of $V_{i-1}$ for $i\geq 1$. When $V=\mathbb{C}[x_{1},\ldots,x_{N}]$ and $V_{i}$ is the set of elements of $V$ that are homogeneous of degree $i$ for all $i\geq 0$, we have that $\partial_{i}$, $\mathcal{D}(\mathcal{R}(\theta))_{i}$, and $d_{i}$ for $i\in[N]$ are examples of such an operator. In this section, we introduce a similar setting that is motivated by the Dunkl bilinear form $\cdot,\cdot]_{\mathcal{R}(\theta)}$ and we continue to discuss this setting in Sections˜4 and 5. See Example˜5.18 for the application of the framework that we introduce to the context of Dunkl operators.
@@ 4. Invertible graded rings of operators
The key contribution that we discuss in this section while proving Theorem˜4.5 is the analysis of the matrices $\{M^{i;L}\}_{i\geq 0}$. This method allows for straightforward proofs that are applicable in a general setting.
@@
    For all $i\geq 0$, there does not exist nonzero $f\in V_{i}$ such that $L(g)f=0$ for all $g\in R_{i}$.
@@
    For all $i\geq 0$, there does not exist nonzero $g\in R_{i}$ such that $L(g)f=0$ for all $f\in V_{i}$.
@@
    The equation $\sum_{j=1}^{|A_{i}|}(L(b_{ij})f)a_{ij}=0$ has no nonzero solutions $f\in V_{i}$ for all $i\geq 0$.
@@ 5.2. Equivariance with respect to a group action and examples
The paper [dunkl_singular_poly] shows that in the setting of Dunkl operators, the invertibility of $L_{(R,V)}$ implies the invertibility of $L$. We repeat the argument from Remark 4.2 of that paper in a more general setting to prove the following result.
@@
For all $h\in H$, $hg^{\prime}=\prod_{h^{\prime}\in H}hh^{\prime}g=g^{\prime}$ by the commutativity of $S$, so $g^{\prime}\in R$.
@@
We can express $g^{\prime}$ as a unique linear combination of elements of $R_{j}$ for $j\geq 0$. Let $m$ be the smallest nonnegative integer $j$ such that the component of $R_{j}$ in this linear combination is nonzero. Then, $m\geq i\geq 1$ because $g\in S_{i}$ and $g^{\prime}\not=0$.
Suppose $f\in V_{m}$. Then, $L(g^{\prime})f\in K$. Suppose $f^{\prime}=\prod_{h^{\prime}\in H\backslash\{1\}}L(h^{\prime}g)f$. Then, $L(g)f^{\prime}\in K$, so if $f^{\prime\prime}$ is the projection of $f^{\prime}$ onto $W_{0}\oplus\cdots\oplus W_{i}$, we have that $L(g^{\prime})f=L(g)f^{\prime}=L(g)f^{\prime\prime}$. However, $L(g)f^{\prime\prime}=0$, which is a contradiction to the invertibility of $L_{(R,V)}$. ∎
@@
Suppose $N\geq 1$, $S=W=K[x_{1},\ldots,x_{N}]$, and that $S_{i}=W_{i}$ is the group of polynomials over $K$ that are homogeneous of degree $i$ for $i\geq 0$. From Lemma˜5.14, any $L\in\mathcal{T}$ is uniquely determined by $L(x_{i})$ for $1\leq i\leq N$. Let $H$ be a finite group which acts on $K^{N}$ such that for all $h\in H$ and $f\in S_{1}$, $\{x\mapsto f(hx)\}\in S_{1}$.
@@
Assume that $L\in\mathcal{H}$ is invertible and that $\text{char}(K)$ does not divide $|H|$. Then, all conclusions of Corollary˜5.16 are true (this includes the conclusions over both $(R,V)$ and $(S,W)$).
@@
We describe the framework that we have developed in the context of Dunkl operators and prove a few well-known results. Consider the setting of Example˜5.13, but set $K=\mathbb{C}$ and $H$ to be $H(\mathcal{R})$, where $\mathcal{R}\subset\mathbb{R}^{N}$ is a finite root system.
Suppose $u_{i}\in\mathbb{R}^{N}$ for $1\leq i\leq N$. Then, we can define $L\in\mathcal{T}$ by $L(x_{i})\triangleq\mathcal{D}_{u_{i}}(\mathcal{R}(\theta))$ after using Lemma˜5.14 and the commutativity of the Dunkl operators. As a special case, we have that $\mathcal{D}(\mathcal{R}(\theta))\in\mathcal{T}$, where we recall that $\mathcal{D}(\mathcal{R}(\theta))(x_{i})=\mathcal{D}_{i}(\mathcal{R}(\theta))$ for $i\in[N]$. For $i\geq 0$, the matrix $M^{i;\,\mathcal{D}(\mathcal{R}(\theta))}$ has entry $(r,s)$ equal to $r,s]_{\mathcal{R}(\theta)}$ for $r\in D_{i}$ and $s\in C_{i}$.
@@
Observe that $E_{a}^{\mathcal{R}(\theta)}(x)$ is the $\Psi$-eigenvector of $\mathcal{D}(\mathcal{R}(\theta))$ and $J_{a}^{\mathcal{R}(\theta)}(x)$ is the $\Psi$-eigenvector of $\mathcal{D}_{H}(\mathcal{R}(\theta))$ for $\Psi:r\mapsto r(a)$, where $a\in\mathbb{C}^{N}$ is fixed. Then, part (B) of Corollary˜5.16 implies that these functions are unique eigenvectors, which is included in the statements of Theorems˜2.5 and 2.8.
The intertwining operator for $(\mathcal{D}(\mathcal{R}(\theta)),\partial)$ is well-studied, where $\partial\in\mathcal{T}$ satisfies $\partial(x_{i})=\partial_{i}$ for $i\in[N]$; equivalently, $\partial\triangleq\mathcal{D}(\mathcal{R}(0))$. As an example of an application of Theorem˜4.5, the theorem implies the well-known result that $(\mathcal{D}(\mathcal{R}(\theta)),\partial)$ intertwines with a unique operator if and only if $\mathcal{D}(\mathcal{R}(\theta))$ is invertible. As explained in [rosler_positivity], in terms of Theorem˜1.3, the intertwiner is given by
$Vf(a)\triangleq\int_{\mathbb{R}^{N}}f(\epsilon)d\nu^{\mathcal{R}(\theta)}_{a}(\epsilon)$
for $f\in\mathbb{C}[x_{1},\ldots,x_{N}]$ and $a\in\mathbb{R}^{N}$. Since the Bessel function is
$J_{a}^{\mathcal{R}(\theta)}(x)=\frac{1}{|H(\mathcal{R})|}V\sum_{g\in H(\mathcal{R})}e^{\left\langle gx,\cdot\right\rangle}(a),$
the characteristic function of $\frac{1}{|H(\mathcal{R})|}\sum_{g\in H(\mathcal{R})}\nu_{ga}^{\mathcal{R}(\theta)}=\nu_{a}^{\text{sym; }\mathcal{R}(\theta)}$ is given by $J_{a}^{\mathcal{R}(\theta)}(ix)$.
@@
For $\theta\geq 0$ and even $i\geq 0$, it is straightforward to compute the decomposition of $M^{i;\,\mathcal{D}_{H}(D^{N}(\theta))}$ when we restrict the rows and columns to $p_{\lambda}$ for $\lambda\in\Gamma_{\text{even};\,N}[i]$ by using the decomposition of $M^{i;\,\mathcal{D}_{H}(BC^{N}(\theta,0))}$. By (20), the decomposition for the remaining entries is given by that of $M^{i;\,\mathcal{D}_{H}(BC^{N}(\theta,1))}$. Furthermore, another interesting question is, what are the decompositions of $M^{i;\,\mathcal{D}(BC^{N}(\theta_{0},\theta_{1}))}$ for $i\geq 0$? Recall that $M^{i;\,\mathcal{D}(D^{N}(\theta))}=M^{i;\,\mathcal{D}(BC^{N}(\theta,0))}$.
@@ 6. Leading order terms for the type A Dunkl bilinear form
6.2. Proof of part (A) of Theorem˜1.2
@@
The next theorem generalizes part (A) of Theorem˜1.2. As we mentioned previously, we focus on the less general version stated in Corollary˜6.11 because it is more relevant for analyzing the convergence of Bessel functions.
@@
6.3. Leading order terms of the type A Dunkl bilinear form in the high temperature regime
Next, we compute the leading order terms of the matrix $([p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}\\ )_{\lambda,\nu\in\Gamma}$ in the regime $\theta N\rightarrow c\in\mathbb{C}$.